Question 8 8 out of 8 points The registrar at Fort Lewis College suspects that business school students are missing too muchSOLVE #10 please!
Answer #1

x

x2

3

9

7

49

2

4

1

1

0

0

2

4

Total

15

67

n=6

Sample mean = 15 EL Σ f

= 2.5

The standard deviation of sample is given by

Σχ2 (Σx) n S = η –1

67- (15) 6 S= 6-1

67 – 37.5 S = 5

S = 5.9

S = 2.43

The 99% confidence interval of mean is given by

x+ta/2, -1) * se(x)

s i t 0.995,5) vn

2.43 2.5 + 4.032143 16

2.5 +3.998

Lower bound= -1.498

Upper bound = 6.498

Answer(8):

Upper bound = 6.498

Answer(9):

Lower bound= -1.498

Answer(10):

We have obtained the 99% confidence interval for the mean number of days absent which is (-1.498, 6.498).

The registrar suspects on average, students are missing about three classes per semester. In the above confidence interval the value 3 is included which indicates that the 3 days is not significantly different from the mean number of days absent.

Hence we can say that the registrar is right about the mean number of days absent with 99% confidence and we can agree with registrar.